OpenStudy (anonymous):

Convert the mixed number, 1 3/4 , to a fraction. Reduce fraction if necessary

OpenStudy (johnweldon1993):

to convert a mixed number \[Whole \frac{ Numerator }{ Denominator }\] You take the denominator (4) in this case....and multiply it by the Whole number (1) in this case...then to this number...you add the numerator (3) in this case So 4 times 1 = 4....add 3 = 7 This number (7) gets put over the original denominator (4) so you have \[1\frac{ 3 }{ 4 } = \frac{ 7 }{ 4 }\]

OpenStudy (anonymous):

yea thanks i just did it to

OpenStudy (johnweldon1993):

no problem